### v2

# MambaFlow: A Mamba-Centric Architecture for End-to-End Optical Flow Estimation

Abstract
Recently, the Mamba architecture has demonstrated significant successes in various computer vision tasks, such as classification and segmentation. However, its application to optical flow estimation remains unexplored. In this paper, we introduce MambaFlow, a novel framework designed to leverage the high accuracy and efficiency of the Mamba architecture for capturing locally correlated features while preserving global information in end-to-end optical flow estimation. To our knowledge, MambaFlow is the first architecture centered around the Mamba design tailored specifically for optical flow estimation. It comprises two key components: (1) PolyMamba, which optimizes feature representation; and (2) PulseMamba, which facilitates efficient flow information dissemination. Our extensive experiments demonstrate that MambaFlow achieves remarkable results. On the Sintel benchmark, MambaFlow records an endpoint error (EPE) of 1.43 and an inference speed of 0.113 seconds, surpassing the state-of-the-art methods including GMFlow (with 18.9% lower EPE and 18.1% faster inference), SeparableFlow (5% lower EPE and 50.5% faster), CRAFT (1.11% lower EPE and 76.5% faster), and DIP (0.7% lower EPE and 77.2% faster)—demonstrating stronger potential for real-world deployment on resource-constrained devices. The source code will be made publicly available upon acceptance of the paper.
1 Introduction
Optical flow estimation, a fundamental problem in computer vision, aims to compute the motion vectors of pixels between consecutive frames in a video sequence. It holds significant and widespread applications in areas such as action recognition [65, 56], video interpolation [16, 51, 28], autonomous driving [59], etc.
Architectures based on Convolutional Neural Networks (CNNs) have achieved pioneering progress in the field of optical flow. FlowNet [5], as the first end-to-end CNNs-based flow network, revealed the immense potential of this learning paradigm. However, it faced limitations in handling large displacements and complex motions. Subsequently, PWC-Net [44] improved estimation accuracy in complex motion scenarios by incorporating pyramid feature extraction and a cost-volume construction, which effectively captured multi-scale local features. Nevertheless, it still encountered challenges in large displacement scenarios. To address these issues, RAFT [45] introduced a recurrent iterative refinement mechanism, which significantly enhanced the accuracy of optical flow estimation. However, such innovation brought a new challenge, which its optimization process relies on numerous iterations, leading to increased parameter counts and longer inference time. Many subsequent variants of the RAFT algorithm have attempted upgrates, such as MS-RAFT [19] and SEA-RAFT [49], but their ability to handle large displacements and complex motions still requires further improvement. To tackle the aforementioned issues, the Transformer architecture, known for its powerful global modeling capabilities, has been introduced into optical flow estimation tasks [15, 42, 30, 43]. Most of them leverage attention mechanism to enhance feature representation, which capture long-range dependencies to significantly improve the accuracy. Notably, unlike optical flow algorithms based on RNNs for flow decoding [45, 20], GMFlow [54] utilizes the attention mechanism to iteratively decode the final optical flow, resulting in faster flow refinement. However, they encounter the quadratic computational complexity of Transformers both during the training and inference stages [54, 43, 15, 42, 30], which severely limits their efficiency in practical applications.
Figure 1: End-point-error on Sintel (clean) vs. inference time (ms) and model params (M). All models were trained on FlyingChairs (C), FlyingThings3D (T), Sintel (S), KITTI (K), and HD1K (H) datasets, with testing performed using one NVIDIA GeForce RTX 3090 GPU. Compared to state-of-the-art models, our proposed model achieves comparable accuracy while significantly reducing inference time, demonstrating an optimal balance between computational efficiency and performance.
Recently, Mamba [8, 10] has been widely used in various computer vision tasks, such as image classification [7, 57, 58],image segmentation [27, 67, 6], object detection [33, 32, 46], image generation [13, 23], video understanding [21, 24, 52], etc. Those remarkable achievements have demonstrated that Mamba can effectively address the challenges posed by CNNs and Transformers, including the local perception limitations of CNNs and the quadratic computational complexity of Transformers [55, 63]. It benefits from a selective state space mechanism, which can efficiently capture global information and handle complex spatiotemporal dependencies with linear complexity [63]. Specifically, Mamba has been conceptually summarized as being ideally suited for tasks involving long sequences and autoregression [60] due to its inherent RNN mechanism within the SSM [10, 9, 8].
Inspired by Mamba’s capabilities for long sequence modeling and autoregression with linear complexity, we propose a novel optical flow estimation framework, MambaFlow, designed to address the challenges in CNNs or Transformer-based methods. As demonstrated in Fig. 1, our method significantly reduces inference time while maintaining remarkblely high accuracy.The framework introduces two core components: (1) the PolyMamba module, which utilizes Self-Mamba, Cross-Mamba, and Multi-Layer Perceptron (MLP) to globally optimize feature representations with lightweight computation, and (2) the PulseMamba module, which effectively combines motion features with Mamba’s autoregressive ability to iteratively generate flow information, addressing challenges such as large displacement and occluded regions.
Specifically, we begins with a CNN backbone to extract robust features from the input image pairs similar to traditional practice [49, 54, 20]. Subsequently, we leverage the suggested PolyMamba to enhance the extracted features, allowing for the capture of complex global information and long-range dependencies. Building on the enhanced features, we further utilize a global matching mechanism [54, 64] to compute similarities for initial flow. At last, we employ PulseMamba to recurrently refine and yield the final precise flow [54]. Our algorithm uses Mamba as the core components in both the encoding and decoding stage, making it an end-to-end optical flow estimation algorithm centered around Mamba. By leveraging Mamba’s capabilities in global modeling and autoregressive properties, as well as its computationally efficient characteristics, the proposed approach achieves a good trade-off between accuracy and performance.
Our main contributions can be summarized as follows:
  • We propose MambaFlow, a novel optical flow estimation architecture built on the Mamba framework both in feature enhancement and flow propagation. To our knowledge, this is the first Mamba-centric architecture designed for end-to-end optical flow estimation. Our method achieves accuracy comparable to state-of-the-art algorithms while significantly reducing training and inference time, which makes it more suitable for real-world application on resource-constrained devices.
  • We introduce the PolyMamba module, which consists of a Self-Mamba, a Cross-Mamba, and a MLP in sequence, to globally enhance feature representation for both the reference image and its cross features with the matching image. As we will demonstrate, it effectively leverages the Mamba’s long sequence modeling capabilities while maintaining a linear computational load.
  • We design the PulseMamba, which includes a motion encoder and a Mamba to facilitate flow propagation, addressing the computational overload, large displacement and occulusion issues in recurrently refined RNN-like methods. By leveraging the Mamba’s powerful autoregressive capability and linear complexity, it offers significant potential to enhance accuracy and speed simultaneously through ablation studies.
2 Related Work
2.1 Optical Flow Estimation
Optical flow estimation was traditionally formulated as an energy minimization task [1, 12, 38], with early variational methods like Horn and Schunck [12] relying on brightness constancy and spatial smoothness assumptions. Subsequent optimization-based approaches improved robustness through regularization [2].
In the era of deep learning, following the introduction of FlowNet [5] as the first CNN-based optical flow method, numerous subsequent algorithms such as PWC-Net [44] and SpyNet [40], adopted a coarse-to-fine strategy for optical flow estimation. However, these approaches struggled with fast motions and small displacements due to inaccurate coarse-level guidance. RAFT [45] addressed this via iterative full-field refinement, inspiring variants such as GMA [20], SeparableFlow [62],and MS-RAFT [19], which introduced modules to better handle occlusion and motion ambiguity. Despite their accuracy, these CNN-based methods require heavy iterative computation and struggle with large displacements and complex motion, limiting their efficiency on resource-constrained devices.
Transformers have emerged as powerful alternatives due to their ability to model global dependencies via self-attention. Models like FlowFormer [15], CRAFT [43], and GAFlow [31] leveraged Transformer architectures to enhance feature expressiveness and global matching. GMFlow [54] and TransFlow [30] further improved efficiency by reformulating optical flow as a one-shot global matching problem. However, the quadratic complexity of attention mechanisms results in high computational and memory costs, limiting deployment on lightweight systems.
2.2 State Space Models
State space models [11], which originated from control theory, have been effectively integrated with deep learning to capture long-range dependencies. Mamba [8] introduced an input adaptation mechanism to enhance the state space model, providing faster inference speed, higher throughput and overall indicators compared to the Transformers of the same size, especially in natural language. Related work on Mamba [50] has shown that it has comparable performance to the Transformer-based methods while significantly reducing computational complexity. This superiority led to further exploration like Vision Mamba [68] and Vmamba [29], which applied SSMs to visual tasks such as image classification [29, 68], video understanding [25, 47], biomedical image segmentation [34, 53] and others [14, 17, 37, 39, 61]. However, their applicability to optical flow estimation has remained unexplored until recently. Lin et al. [26] pioneered the use of Mamba in the scene flow estimation task, proposing an SSM-based iterative update module to replace the GRU module as a decoder. Nevertheless, it was specifically designed for scene flow with point cloud feature and is not suitable for the generally used backbone optical flow methods.
Inspired by the Mamba’s strength of long sequence modeling and autoregression with linear complexity, we propose a Mamba-centric architecture for optical flow estimation. Our goal is to design an end-to-end solution that utilizes Mamba for feature enhancement and flow optimization. During the feature enhancement phase, PolyMamba enhances features from pairs of images by sequentially employing Self-Mamba, Cross-Mamba, and MLP to capture global information and mutual dependencies. At the flow refinement phase, PulseMamba iteratively optimize the fina optical flow by leveraging a motion encoder and a SSM. These Mamba-centric designs guarantee the accuracy of our method while maintaining its high speed in training and inference, providing valuable insights for resource-constrained devices.
Figure 2: The overall network architecture of our MambaFlow.
3 Method
3.1 Preliminaries
SSMs are inspired by linear systems, mapping an input sequence $x_{t}\in\mathbb{R}$ to an output $y_{t}\in\mathbb{R}$ through a hidden state $h_{t}\in\mathbb{R}^{N}$. The model can be formulated as linear ordinary differential equations (ODEs):
$h_{t}=\overline{\mathbf{A}}h_{t-1}+\overline{\mathbf{B}}x_{t},$ (1)
$y_{t}={\mathbf{C}}h_{t},$ (2)
where $\overline{\mathbf{A}}$ and $\overline{\mathbf{B}}$ represent the discretized parameters derived from the given parameters $\mathbf{A}\in\mathbb{R}^{N\times N}$, $\mathbf{B}\in\mathbb{R}^{N\times 1}$, and the time scale parameter $\boldsymbol{\Delta}$:
$\overline{\mathbf{A}}=\exp(\boldsymbol{\Delta}\mathbf{A}),$ (3)
$\overline{\mathbf{B}}=(\boldsymbol{\Delta}\mathbf{A})^{-1}(\exp(\boldsymbol{\Delta}\mathbf{A})-\mathbf{I})\cdot\boldsymbol{\Delta}\mathbf{B}.$ (4)
Additionally, $\mathbf{C}\in\mathbb{R}^{1\times N}$ serves as the projection matrix mapping the hidden state $h_{t}$ to the output $y_{t}$.
Finally, the models generate the output using a global convolution operation, defined as:
$\overline{\mathbf{K}}=(\mathbf{C}\overline{\mathbf{B}},\mathbf{C}\overline{\mathbf{A}}\overline{\mathbf{B}},\ldots,\mathbf{C}\overline{\mathbf{A}}^{M-1}\overline{\mathbf{B}}),$ (5)
$y=x\ast\mathbf{\overline{K}},$ (6)
where $M$ represents the length of the input sequence $x$, and $\mathbf{\overline{K}}\in\mathbb{R}^{M}$ denotes a structured convolutional kernel.
The recently proposed advanced state-space model, Mamba, enhances $\mathbf{B}$, $\mathbf{C}$, and $\boldsymbol{\Delta}$ by making them input-dependent, thereby enabling dynamic feature representations. Specifically, these matrices are defined as follows:
$\mathbf{B}_{i}=S_{B}(x_{i}),$ (7)
$\mathbf{C}_{i}=S_{C}(x_{i}),$ (8)
$\boldsymbol{\Delta}_{i}=\text{Softplus}(S_{\Delta}(x_{i})),$ (9)
where $S_{B}$, $S_{C}$, and $S_{\Delta}$ are linear projection functions, and $i$ denotes the index of the $i$-th element in the sequence.
Simultaneously, the parallel scan algorithm enables Mamba to leverage the same parallel processing advantages as described in (5), thereby facilitating efficient training.
3.2 Overall Architecture
Optical flow can be intuitively understood as a task aimed at finding pixel-wise correspondences between consecutive images. As shown in Fig. 2, our proposed framework, MambaFlow, is centered around two novel modules: the feature enhancement module PolyMamba and the flow refinement module PulseMamba.
Given two input frames $\boldsymbol{I}_{1}$ and $\boldsymbol{I}_{2}$, we first extract dense feature maps $\boldsymbol{F}_{1},\boldsymbol{F}_{2}\in\mathbb{R}^{H\times W\times D}$ using a shared-weight convolutional encoder, where $H$, $W$, and $D$ denote the height, width, and feature dimension of the features. These features are then processed by the PolyMamba module, which stacks Self-Mamba and Cross-Mamba blocks to capture both intra-frame and inter-frame dependencies. To further enhance representation capacity, a lightweight MLP layer is appended after each block group for local-global feature fusion. The resulting features $\boldsymbol{\mathit{F}}_{q}$ and $\boldsymbol{\mathit{F}}_{v}\in\mathbb{R}^{H\times W\times D}$ are then used for global matching to estimate the initial flow.
The initial prediction $\mathbf{V}_{\text{initial}}$ is then refined by the PulseMamba module, which incorporates cost volume and motion features into a stack of Mamba layers for iterative flow refinement. This module significantly improves performance in occluded or textureless regions, enhancing both the accuracy and robustness of optical flow estimation.
3.3 Self-Mamba
The input features $\boldsymbol{\mathit{F}}_{1},\boldsymbol{\mathit{F}}_{2}\in\mathbb{R}^{H\times W\times D}$ are enriched with learnable positional embeddings to retain spatial positional information. Subsequently, $\boldsymbol{\mathit{F}}_{1}$ and $\boldsymbol{\mathit{F}}_{2}$ are processed through a weight-sharing Self-Mamba block, yielding enhanced representations denoted as $\boldsymbol{\mathit{F}}_{1}^{\text{self}}$ and $\boldsymbol{\mathit{F}}_{2}^{\text{self}}$, respectively.
The processing within each Self-Mamba block can be summarized as a bidirectional operation, where the forward and backward hidden states are computed and combined to enrich the feature representation. The overall process is expressed as:
[…]
Here, the Self-Mamba function represents the state space model used to compute the forward and backward hidden states, with the sequence $\boldsymbol{\mathit{F}_{i}}$ being reversed to distinguish between the forward and backward directions.
3.4 Cross-Mamba
While the Self-Mamba architecture is effective in modeling long-range dependencies within a single sequence, it lacks the capacity to capture inter-sequence relationships—an ability that is crucial in optical flow estimation, where information from two correlated feature maps must be jointly reasoned. To address this limitation, we introduce a cross-sequence interaction module based on selective state-space modeling, enabling rich bidirectional information exchange between paired features.
Given the enhanced features $\boldsymbol{\mathit{F}}_{1}^{\text{self}},\boldsymbol{\mathit{F}}_{2}^{\text{self}}$, the main branch applies a depth-wise causal convolution and activation:
$\boldsymbol{x}_{\mathrm{conv}}=\texttt{SiLU}\left(\texttt{DWConv}(\boldsymbol{\mathit{F}}_{1}^{\text{self}})\right),$ (11)
while the modulation branch linearly projects the second stream:
$\boldsymbol{x}_{\mathrm{mod}}=\texttt{Proj}(\boldsymbol{\mathit{F}}_{2}^{\text{self}}).$ (12)
Their concatenation is projected into three parameter groups:
$[\,\boldsymbol{\Delta},\;\mathbf{B},\;\mathbf{C}\,]=\texttt{Linear}([\boldsymbol{x}_{\mathrm{conv}};\;\boldsymbol{x}_{\mathrm{mod}}]),$ (13)
The hidden state $\boldsymbol{y}$ is then updated through a SelectiveScan kernel, which dynamically integrates temporal and spatial information using parameterized state-space operations.
A final linear projection yields the cross-attended representation:
$\boldsymbol{\mathit{F}}^{\text{cross}}=\texttt{OutProj}(\boldsymbol{y}).$ (14)
To model temporal symmetry, Cross-Mamba is applied in both forward and backward directions—just as in the Self-Mamba block—and the outputs are summed:
[…]
3.5 Global Matching
After extracting the enhanced features $\boldsymbol{\mathit{F}}_{q}$ and $\boldsymbol{\mathit{F}}_{v}$ from our PolyMamba module, we construct a 4D cost volume of size $H\times W\times H\times W$ by computing the dot-product similarity between all pixel pairs from both features. The constructed 4D cost volume can be interpreted as a collection of $H\times W$ 2D cost maps, where each map evaluates the feature similarity between a single pixel in $\boldsymbol{\mathit{F}}_{q}$ and all pixels in $\boldsymbol{\mathit{F}}_{v}$. To obtain a matching distribution, a softmax operation is applied across the last two dimensions of the cost volume. The construction process is expressed as:
$\mathbf{M}=\text{softmax}(\mathbf{\frac{\boldsymbol{\mathit{F}}_{\mathit{q}}\,\boldsymbol{\mathit{F}}_{\mathit{v}}^{\top}}{\sqrt{D}}}).$ (16)
where $D$ is the feature dimension. To estimate the optical flow, the correspondence is obtained by taking a weighted average of the 2D pixel grid coordinates $\mathbf{G}\in\mathbb{R}^{H\times W\times 2}$ in the target image with the matching distribution $M$. Finally, the optical flow $\mathbf{V}_{\text{initial}}$ is calculated as the difference between the corresponding pixel coordinates:
$\mathbf{V}_{\text{initial}}=\mathbf{M}\mathbf{G}-\mathbf{G}.$ (17)
The global matching method effectively improves the accuracy and stability of optical flow estimation by capturing global contextual information.
3.6 PulseMamba
To effectively address the challenges of occlusions and boundary conditions in optical flow estimation, we propose a novel module named PulseMamba. PulseMamba integrates cost volume information and motion features, achieving dynamic and context-aware optical flow propagation through iterative refinement.
Given the initial optical flow $\mathbf{V}_{\text{initial}}$, the source image features $\boldsymbol{F}_{q}$, and the 4D cost volume (denoted as $\mathbf{C}$), PulseMamba proceeds as follows:
First, motion features are extracted by encoding the initial flow and the 4D cost volume:
$\boldsymbol{M}=\text{MotionEncoder}(\mathbf{V}_{\text{initial}},\mathbf{C}),$ (18)
PulseMamba employs multiple stacked Mamba layers for iterative feature refinement. At each iteration $i$, the input is formed by concatenating the image features, motion features, and current flow prediction, and the flow is updated as:
$\mathbf{V}^{(i)}=\text{Mamba}^{(i)}\left([\boldsymbol{F}_{q};\;\boldsymbol{M};\;\mathbf{V}^{(i-1)}]\right),$ (19)
where $\mathbf{V}^{(0)}=\mathbf{V}_{\text{initial}}$.
The output flow from each iteration becomes the input for subsequent refinement. After $N$ iterations, the final refined optical flow is obtained:
$\mathbf{V}_{\text{final}}=\mathbf{V}^{(N)}.$ (20)
Through this iterative and context-aware refinement, PulseMamba significantly improves the accuracy and robustness of optical flow estimation, particularly under complex conditions involving occluded regions and ambiguous boundaries.
Table 1: Quantitative results on Sintel and KITTI. MambaFlow achieves at least 18% faster inference on the KITTI (test) dataset compared to other state-of-the-art methods with competitive performance. All compared methods were evaluated under the same training configuration, and a unified "C+T+S+K+H" training strategy was adopted to ensure the fairness of the results. For performance measurements, we used the normalized setting of batch size 1 and input resolution 540 × 960. In addition, all inference times are measured on the RTX 3090. The significant performance improvement highlights MambaFlow’s advantages in speed and efficiency. It is worth noting that while maintaining competitive accuracy on the Sintel and KITTI datasets, we have successfully achieved a significant improvement in speed, further verifying the superiority of MambaFlow in the speed-accuracy balance.
Training Data Method Sintel KITTI Inference Cost
[…]
C+T+S+K+H RAFT [45] 1.61 2.86 5.10 5.3M 140.7ms
SeparableFlow [62] 1.50 2.67 4.64 6.0M 229.1ms
GMA [20] 1.39 2.47 5.15 5.9M 183.3ms
DIP [66] 1.44 2.83 4.21 5.4M 498.9ms
GMFlowNet [64] 1.39 2.65 4.79 9.3M 244.3ms
CRAFT [43] 1.45 2.42 7.79 6.3M 483.4ms
MS-RAFT [18] 1.37 2.67 4.88 13.5M 627.3ms
MatchFlow [4] 1.33 2.64 4.72 14.8M 283.8ms
MS-RAFT+ [19] 1.23 2.68 4.15 16.0M 1108.2ms
GMFlow [54] 1.74 2.90 9.32 4.7M 138.5ms
MambaFlow (ours) 1.43 2.45 8.06 11.2M 113.5ms
4 Experiments
Datasets.
Following the previous methodology, we first pre-train on the FlyingChairs (Chairs)[5] and FlyingThings3D (Things)[35] datasets. For fine-tuning, we adopt a mixed training set consisting of KITTI [36], HD1K [22], FlyingThings3D, and Sintel [3] training sets. Finally, we further fine-tune the model on the KITTI training set and report the performance on the respective online benchmarks.
Metrics.
In our experiments, we employed evaluation metrics commonly used in optical flow estimation. Firstly, we utilized the end-point-error (EPE) to measure the average Euclidean distance between the predicted and ground truth optical flow vectors. For the KITTI dataset, we adopted the F1-all metric, which represents the percentage of pixels with an error exceeding 3 pixels or a relative error greater than 5%. To gain deeper insights into the model’s performance under large motion magnitudes, we defined $s_{40+}$ as the EPE for pixels where the ground truth flow magnitude exceeds 40 pixels.
Implementation Details.
Our model is constructed using the PyTorch library and trained on 8 NVIDIA GeForce RTX 3090 GPUs. Following the training strategy from previous works [54, 45], we use the AdamW optimizer. The model is first pre-trained on the Chairs dataset for 100k iterations with a batch size of 16 and a learning rate of $4\times 10^{-4}$. Then, it is fine-tuned on the Things dataset for 200k iterations with a batch size of 8 and a learning rate of $2\times 10^{-4}$. During this process, we select the best-performing model on the Things dataset and fine-tune it for an additional 800k iterations to validate the ablation study. Subsequently, the model fine-tuned for 200k iterations on the Things dataset is used as the initial model and further trained on the Sintel dataset, keeping the same hyperparameters as those used for the Things dataset, including a batch size of 8 and a learning rate of $2\times 10^{-4}$. Finally, we continue training on the KITTI-2015 training set for 200k iterations with the batch size and learning rate unchanged.
Figure 3: Visual results on Sintel test set.
Table 2: PolyMamba components. Ablation study on individual components, revealing that Cross-Mamba contributes the most.
setup Things (val) Sintel (train) Param (M)
[…]
full 2.91 - 1.23 - 2.76 - 11.2 -
w/o self 3.11 $\uparrow$6.91.27 $\uparrow$3.32.82 $\uparrow$2.28.9 $\downarrow$24.6%
w/o cross 5.01 $\uparrow$72.22.08 $\uparrow$69.14.18 $\uparrow$51.45.4 $\downarrow$54.2%
w/o MLP 3.22 $\uparrow$10.71.33 $\uparrow$8.12.96 $\uparrow$7.28.2 $\downarrow$30.5%
w/o pos 3.00 $\uparrow$3.11.26 $\uparrow$2.42.82 $\uparrow$2.211.2 -
Table 3: Number of PolyMamba blocks. Using 8 blocks of Mamba strikes an optimal balance between accuracy and parameter efficiency.
blocks Sintel (train) KITTI (train) Param (M)
[…]
4 1.28 2.94 10.03 29.18 7.2
6 1.24 2.87 9.64 27.23 9.2
8 1.23 2.77 8.99 26.84 11.2
10 1.22 2.82 8.96 26.52 13.2
12 1.23 2.69 8.53 25.70 15.2
24 1.21 2.63 8.38 24.69 27.3
4.1 Comparison with State-of-the-Arts
We compared the performance of different optical flow algorithms on the Sintel and KITTI datasets, and the experimental results are shown in Table 1.
Results on Sintel.
MambaFlow achieved an EPE of 1.43 on the Sintel Clean dataset and an EPE of 2.45 on the Final dataset. Fig. 3 presents a qualitative analysis of optical flow estimation, comparing MambaFlow with other state-of-the-art approaches on the Sintel test set. The results show that MambaFlow’s performance on Clean and Final is at the same level as current mainstream methods such as RAFT and GMA, demonstrating its strong adaptability in both simple and complex scenarios.
Results on KITTI.
MambaFlow achieves an F1-all score of 8.53% on the KITTI dataset while demonstrating a significant advantage in inference speed. Specifically, MambaFlow requires only 113.5ms for inference on an RTX 3090, making it 18.1% faster than GMFlow (138.5ms), which is the fastest among the compared methods. Furthermore, MambaFlow runs 2.5 times faster than MatchFlow (283.8ms) and nearly 10 times faster than MS-RAFT+ (1108.2ms), while maintaining competitive accuracy. This demonstrates that MambaFlow is able to maintain comparable accuracy to mainstream methods while significantly reducing inference time when handling larger and more challenging scenarios.
4.2 Ablation Analysis
Components of PolyMamba.
We ablate different PolyMamba components in Table 2. Among them, Cross-Mamba contributes the most as it models the mutual relationship between two features, a relationship that is otherwise missing in the features extracted from the convolutional backbone. Self-Mamba aggregates contextual cues within the same feature, resulting in an additional performance improvement, and positional information introduces location dependence into the matching process, helping to mitigate ambiguities present in feature-similarity-based matching alone. Removing the MLP reduces the number of parameters, but also results in a slight decrease in performance.
Number of PolyMamba blocks.
We conduct an ablation study on the number of PolyMamba blocks, as shown in Table 3. The experiment compares the impact of varying the number of PolyMamba blocks on overall performance. Among all the tested configurations, the PolyMamba with 8 blocks achieves a good balance between accuracy and parameters.
Number of PulseMamba iterations.
As shown in Table 5, we investigated the impact of different iterations of PulseMamba on EPE on the Sintel dataset. As the number of iterations increases from 1 to 3, the overall EPE gradually decreases from 0.77 to 0.65; the error of the matched area decreases from 0.58 to 0.49, and the error of the unmatched area decreases significantly from 3.68 to 3.16, with the most obvious improvement. This trend shows that more iterations can improve the prediction accuracy of the occluded area.
Weight sharing vs. non-sharing comparison.
We compared Mamba modules with shared and independent weights at the same number of iterations. As shown in Table 5, the non-shared scheme achieved slightly better results on all metrics, matching and non-matching metrics. This is because the parameters of each layer can more flexibly adapt to feature changes between iterations. For example, in two iterations, shared weights reduced the model size by 7.5% (from 12.0 to 11.1), but caused the error to increase by 5.9% (from 0.68 to 0.72). These findings suggest that for complex matching tasks, it is reasonable to accept a moderate increase in the number of parameters due to the improvement in feature representation ability.
Training iterations.
As shown in Table 5, Mambaflow has achieved certain improvements in performance as the number of training iterations increases. This shows that more training iterations can effectively improve the performance of the model, making it more stable and accurate in subsequent tasks.
Table 4: Ablations on PulseMamba designs.
PulseMamba iterations share Mamba? Sintel (train, clean)
all matched unmatched
0 2.97 0.68 15.75
1 0.77 0.58 3.68
[…]
Table 5: Training iterations.
[…]
5 Conclusion
As Mamba continues to gain significant progresses in the field of computer vision, this paper explores for the first time to leverage it to estimate optical flow in an end-to-end paradigm. We suggest a Mamba-centric optical flow method, MambaFlow, which contributes two core Mamba-based components: PolyMamba and PulseMamba, proposed to accurately and efficiently enhance feature representation and direct the ultimate flow, respectively.The proposed approach achieves a good balance between accuracy and performance, comparing to the state-of-the-art methods. We attribute this to the Mamba’s capabilities in global modeling and autoregression, as well as its linear complexity, which offers greater potential for real-world deployment on resource-constrained devices, and provides valuable insights that may influence ongoing and future developments in this rapidly evolving field of optical flow.
Limitations.
While MambaFlow achieves results comparable to or superior to some state-of-the-art methods, there is still room for future improvement, as shown from the KITTI results in Table 1. Initially, it might not generalize well when there exists a substantial disparity between the training data and the test data. Secondly, it can still provide better accuracy under adverse conditions, such as motion blur, fog, and high-resolution, which remain unexplored. We will pretrain our model on large-scale datasets such as TartanAir [48] and VIPER [41], which can escalate its generalization ability. Furthermore, there are numerous outstanding variants of Mamba [63] have emerged recently, enabling us to adjust the entire architecture and its core Mamba-based components for more accurate estimations and to facilitate real-world applications.